OpenStudy (anonymous):

Given the system of equations, what is the solution? x + 2y = 7 x - 2y = -1 {(-8, -12)} {(3, 2)} {(-4, 6)}

OpenStudy (anonymous):

3,2 adding both the eqn we get 2x=6 i.e x=3

OpenStudy (anonymous):

and substitiing the value of x=3 in any of the eqn we get:3+2y=7 2y=4=>y=2

OpenStudy (anonymous):

x+2y=7 (1) x-2y=-1 (2) (1) + (2) => 2x = 6 => x= 3 combined with (1) => 3 + 2y = 7 => y =2. So: (x,y) = {(3,2)} is right!